UT Oct 30 - D/E LBTO Night


Observers: B. Rothberg, M.Wagner, O.Kuhn

OSA: J. Williams

ISp: J. Riedl

Mtn Manager: C. Bolyard

Plan for Night:

1. AO Checkout

2. MODS D-time Science

3. LUCI-1/LUCI-2 AGW Transform confirmation

4. LUCI/MODS D-time science time permitting after transforms

Backup:

1. MODS Pointing Model

2. Poor Weather D-Science


Sunset-UT 02:22 - AO Checkout with Josh and Al - conditions verying clouds

UT 02:26 - reconfiguring for MODS-Binocular

UT 03:17 - acquiring V339Del

UT 03:19 - starting spectroscopy

UT 03:27 - spectra emission lines faint - changing itimes

0005-0006 15sec MODS-2 Blue

0002-0003 15sec MODS-1 Blue

0005-0007 15sec MODS-1 Red

Changed to:

2x360 MODS-2 Blue,

2x360 MODS-1 Red

4x180 MODS-1 Blue - (this was an error, should have been 2x360)

Completed UT 03:46

UT 03:48 - acquiring V809Cep

UT 03:55 - conditions getting worse - clouds thickening

UT 03:57 - can't see it in 60sec acquisition image in Red or Blue - trying 120sec - if can't see target then likely not doable

UT 04:01 - punting target -going to Feige 110

UT 04:09 - starting spectra spectra Feige 110

UT 04:21 - complete Feige 110 3x90dec

UT 04:26 - can't see anything except Mars

UT 04:41 - going back to V809ep to try again

UT 04:46 - found target on SX, DX missed, doing pointing check

UT 04:49 - resending preset

UT 04:59 - red is taking extra long to read out

UT 05:02 - starting sience exposures - first exposures:

mods1b.20181030.0011-0013

mods1r.20181030.0021-0023

mods2b.20181030.0020-0022

UT 05:12 - DIMM ~0.9" Guider WFS 0.59" - more of the sky has opened up - thick clouds to the NE

UT 06:09 - scripts complete

Rothberg IMBH Candidate target

UT 06:12 - switching to LUCI-Bino, moved Left rotator zeropoint back to -187.917 (also for ARGOS) - once new transform is uploaded will set back to -187.315 for both

UT 06:54 - acquiring Tellurc star - can't use LUCI on RM507 machines - too slow and Aladdin won't sync data

UT 06:57 - starting Tellruic spectra - do not use images luci1.20181030.0032/0033 and luci2.20181030.0009/0010 as the central wavelength was set incorrecly.

Manually taking the zJ spec data

luci1.20181030.0034 - +10" zJ 1.17micron 25x2.51sec

luci2.20181030.0011 - +10" zJ 1.17micron 25x2.51sec

luci1.20181030.0035 - -10" zJ 1.17micron 25x2.51sec

luci2.20181030.0012 - -10" zJ 1.17micron 25x2.51sec

-collimation went haywire on rightside during last telluric exposure

UT 07:10 - moving to science target

UT 07:15 - collimation went nuts on both sides

UT 10:40 - closing up - can't see anything

UT 07:26 - putting masks into FPU - going to nearby brght star to collimate

UT 07:39 - mirror panic on SX

UT 07:51 - returning to bright collimation star near target

UT 07:54 - left side panicked again

UT 08:08 - going to LUCI-2 only, SX mirror unavailbale

UT 08:14 -sending LUCI-2 preset to science target

UT 08:22 - I can see the blob in the acq image - puttingit at 1488 on the longslit

UT 08:26 - starting spectroscopy - taking what i can get

UT 08:28 - WFS is a mess

UT 08:34 -Josh attempting to manually adjust zernikes

UT 08:45 - DIMM now 1.3-1.4"

UT 08:56 - finishing exposure #8 (L2 0017-0024) - then will move to zenith to fix issues

UT 09:08 - resending science script

UT 09:09 - GS not found

UT 09:16 - collimation struggling again on return to field, going to zenith to try and recover SX - will then decide what to next

UT 09:19 - LUCI-2 issue with compensation mirror - cannot turn off flexure

UT 09:21 - stopped and restarted compensation mirror on LUCI-2

UT 09:47 - reconfiguirng for MODS-Bino - LUCI AGW transform postponed until Thursday

- thanks to slowness of using OBS2 over the OpenNX, i've lost text from UT 09:50-UT 11:39 due to an errant keystroke

-closed for calsibrations, didn't take any MODS data (attempted)

Cals done for LUCI-2,

UT 11:39 MODS-1 Cals show a problem - QTH lamps don't turn on

mods1r.20181030.0032-0034, mods1b.20181030.0027-0032 QTH no flux

mods1r.20181030.00335-0037 VF10 OK

Taking MODS-2 Blue Biases 8k

Testing QTH1 Lamp MODS-1

mods1b.20181030.0033 QTH1 10 8sec - no flux

mods1b.20181030.0034 QTH2 5 2 sec - counts appear

mods1b.20181030.0035 QTH1 5 5sec - no flux

mods1b.2018103.0036 - turning on QTH1 at 10 wait 30seconds

UT 11:58 - putting MODS-2 to sleep

UT 11:59 - taking MODS-1 Biases 8k

Submitted IT 7461 regarding QTH1 lamps on MODS-1

UT 12:09 - done with MODS cals (except for slitflats) -putting MODS1 to sleep and going home.

-- BarryRothberg - 30 Oct 2018
Topic revision: r3 - 30 Oct 2018, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback